The area of a square is 3 less than the perimeter of the same square. What is the length of one side of the square.
ANY HELP WOULD BE AWESOME...THANKS EVERYONE

Answer by checkley75(3666)

X^2=4X-3
X^2-4X+3=0
(X-3)(X-1)=0
X-3=0
X=3 ANS.
X-1=0
X=1 ANS.
PROOFS:
3^2=12-3
9=9
1^2=4*1-3
1=4-3
1=1